When x = 4, y = 1/2

Work out the value of y when x = 2


1
Expert's answer
2021-01-29T05:03:03-0500

Proportion


"\dfrac{4}{1/2}=\dfrac{2}{a}"

"a=\dfrac{2(1/2)}{4}"

"a=\dfrac{1}{4}"

When x = 2, y = 1/4